# The Loss of Control Playbook: Degrees, Dynamics, and Preparedness

### v1

@@
The Loss of Control Playbook: Degrees, Dynamics, and Preparedness
@@ Introduction
In turn, this allowed us to extrapolate some defining elements of LoC scenarios and to support decision- and policymakers in carving out the content and boundaries of LoC more clearly. Our analysis further enabled us to assess why LoC definitions across the literature may differ and to propose a novel conceptualization of LoC. Specifically, in our literature review, we found that LoC scenarios often differ in terms of their severity and persistence. These two metrics broadly capture intuitions around, among other things, the number of people affected, the degree to which those people are affected, and the difficulty of stopping the LoC incident. We subsequently used these metrics to plot concrete LoC scenarios from the literature (see Figure 2) onto an experimental graph, allowing us to provide additional context to existing descriptions of LoC.
@@ 1.A.1 Findings from Loss of Control Definitions in Literature
Despite certain commonalities, such as concerns about AI systems operating beyond the scope of reliable human direction or oversight, definitions in the broader AI literature differ in their emphasis and framing [bernardi2025societaladaptationadvancedai, kulveit2025gradualdisempowermentsystemicexistential, UK_DSIT_FrontierAI_2023, CIGI_Framework_Convention_2024] [^3] Due to our methodology, as described later in this Chapter, human disempowerment, due to its uncertain nature, is out of the scope of this report.thereby presenting conceptual distinctions. Divergences between definitions can make it difficult to implement functional frameworks and enact suitable interventions. One axis of variation relates to the cognitive capabilities required by an AI system for LoC. Some researchers appear to suggest LoC may occur once AI systems become more intelligent than humans [hendrycks2023overviewcatastrophicairisks, MIRI_Mission_Strategy_Update_2024]. Others more explicitly refer to artificial superintelligence when describing LoC risk [RAND_Rise_Fall_Nations, barnett2025aigovernanceavoidextinction]. In contrast, Stuart Russell’s definition in ‘Artificial Intelligence and the Problem of Control’ requires only a “sufficiently capable machine” [Russell2022], rather than one explicitly surpassing human intelligence or achieving superintelligence. A further divergence across definitions involves the difference in treatment between oversight and control. While some definitions refer to control [CA_Frontier_AI_Policy_2025, IDAIS_Shanghai_2025], others emphasize oversight instead, focusing on AI systems that “operat[e] beyond human oversight” [METR_AGI_Definitions_2025]. Another axis of variation concerns the possibility of regaining control. For example, the COP [EU_AI_COP] does not necessarily imply that LoC is irreversible, whereas the IASR [IASR_2025], the Singapore Consensus [bengio2025singaporeconsensusglobalai], and other reports [Bengio_managing_risks_2024, hendrycks2023overviewcatastrophicairisks] all suggest that LoC involves the absence of a clear path to regaining control. These interpretations imply that once control is lost, recovery may be impossible or extremely difficult. Overall, we found that while definitions aim to cover similar areas, they do so in sufficiently diverse ways to prevent a common, clear definition of LoC from arising from the AI literature.
@@ 1.A.2 Findings from Loss of Control Scenarios in Literature
In total, we found 12 text passages that fulfilled all of our criteria and were therefore classified as concrete LoC scenarios. See Figure 1 for a high-level oversight of how our methodology worked.
@@
Next, we plotted all 12 concrete LoC scenarios on an experimental graph, using severity and persistence as the axes, [^13] Given the difficulty of computing the precise economic impact for persistence, especially for concrete LoC scenarios from the literature that do not tend to consider this aspect in any detail, we decided to use the same economic impact estimate for persistence as for severity. We note that, while severity and persistence are not necessarily equivalent in terms of economic impact in all instances, we believe that they are quasi-equivalent for the scenarios plotted on our graph (see Figure 2). and color-coded them within their respective threat categories (see Figure 2). In doing so, we noted that certain areas emerged from our plots and sought to contextualize them within concepts and frameworks that already exist in governance and which decision- and policymakers may be familiar with.
@@
First, we sought to contextualize the area indicated in the lower left of our graph (see Figure 2). In doing so, we came across the economic consequence thresholds that the U.S. Department of Homeland Security (DHS), alongside federal agencies and the intelligence community, identified as “necessary to create a national level-event” in the Strategic National Risk Assessment (SNRA) [DHS_SNRA_2011]. Specifically, in 2011, DHS led an effort to “identify the types of incidents that pose the greatest threat to the Nation’s homeland security,” including various natural, technological, and adversarial, human-caused hazards [DHS_SNRA_2011]. [^16] We decided to adopt the SNRA as a term of reference because the U.S. is the default location where the majority of literature assumes LoC scenarios to occur. For instance, ‘Scenario 1’ in [Kalra_GeopoliticsAGI] is an example of a scenario that envisages AI-caused electricity blackouts starting in the U.S. In some cases, the location of the LoC scenario described by literature is ambiguous (for instance, ‘Scenario 4’ in [Kalra_GeopoliticsAGI]), so for the purposes of making a comparable economic impact estimate, we assumed that the scenario occurred within the U.S. to ensure consistency across scenarios. For example, earthquakes, floods, hurricanes, wildfires, and cyberattacks against physical infrastructure meet the threshold if they result in direct economic losses exceeding \$100 million [DHS_SNRA_2011]. Similarly, a cyberattack against data  [^17] In this context, data refers to the information contained in a computer system or data processes. is considered a national-level event if it results in economic losses of \$1 billion or more. In this research report, we adopt this last threshold set by the SNRA at \$1 billion or greater (adjusting it for inflation to approximately \$1.4 billion) to provide some contextualization of LoC scenarios (see Figure 2). We refer to this threshold, i.e., the \$1.4 billion, for convenience, as the ‘national risk assessment threshold.’ In summary, the national risk assessment threshold is the highest threshold based on economic consequences that DHS, federal agencies, and the intelligence community identified in the Strategic National Risk Assessment to demarcate a national-level event (i.e., a threat or hazard that has the potential to significantly impact the U.S. homeland security). We indicate this threshold in orange in Figure 2.
@@
Second, we sought to contextualize the area indicated in the top right of our graph (see Figure 2). In doing so, we drew on existing conceptualizations describing events of absolute scale and permanence. In other words, events that could lead to the destruction of humanity’s long-term potential [ord2020precipice, sundaram_mani_2025, stauffer2023existential, DallasFed_AI_Productivity_2025]. [^18] We note that we were unable to find a numerical operationalization for ‘existential catastrophe’ in the literature. Therefore, the boundary given on this graph is for illustrative purposes only, but is likely accurate, as it captures an absolute scale and permanence akin to the best interpretation of an existential catastrophe (see Figure 2). These events are commonly referred to as an ‘existential catastrophe,’ and we adopted this term to establish our upper boundary. In summary, existential catastrophe demarcates the point at which humanity loses control over its future in an absolute sense. We indicate this threshold in bold red dashes in Figure 2.
@@
  • Deviation: captures events that cause some harm or inconvenience, but lacks the requisite severity and persistence for inclusion within the threats and hazards that have the potential to significantly impact national preparedness. The events in this category would be plotted in the lower left corner of our graph (see Figure 2), and delineated from LoC scenarios in other categories through the national risk assessment threshold.
  • Bounded LoC: capture events that cause great damage or suffering, and are difficult, but possible, to contain, albeit potentially at great cost. The events in this category are mapped between the lower left corner and the top right corner.
  • Strict LoC: captures events which are maximally severe and permanent, such as events that result in humanity as a whole becoming extinct. The events in this category are mapped at the top right corner of our graph (see Figure 2).
@@ 1.B.1 Deviation
This category captures scenarios where an AI system briefly deviates from human intent, and, while causing some real-world damage or minor economic impacts, remains below the threshold for inclusion within the threats and hazards that have the potential to significantly impact the U.S. homeland security based on the Strategic National Risk Assessment [DHS_SNRA_2011]. These deviations from human intent are simple to stop at low or no cost, meaning that scenarios in this category would be plotted in the lower left corner on our graph (see Figure 2), and are delineated from other categories by what we term the national risk assessment threshold. Overall, we propose that the national risk assessment threshold is a principled boundary since it is calibrated to exclude events where harm is limited and containment is easy, that is, deviations.
@@
Nevertheless, we offer our reasoning as to why we do not believe the COP definition [EU_AI_COP] should be interpreted so broadly as to include the category of Deviation, derived from our earlier review and methodology. A broad interpretation of “losing the ability to reliably direct” would already capture scenarios that occur today. Consider the following two examples. In one instance, an AI system recently deleted an entire database, even though it was instructed not to modify the code [Cybernews_Replit_VibeCoding_2025]. In another instance, OpenAI’s agent Operator conducted a purchase without requesting the users’ consent, despite having been told only to find the cheapest option [Fowler_WaPo_Operator_2025]. This occurred even though the agent was trained to ask the humans’ consent before conducting irreversible actions such as purchasing goods or sending emails [OpenAI_introducing_operator]. In both cases, humans were not able to “reliably direct” an AI system and, therefore, both cases could, in theory, qualify as LoC based on a broad interpretation of “losing the ability to reliably direct.” However, we propose that it is unlikely the LoC definition in the COP was supposed to be interpreted so extensively as to include similar cases, which we believe qualify as Deviation rather than LoC. First, including the category of Deviation within the definition of LoC in the COP [EU_AI_COP] would run the risk of making the LoC definition overly broad, rather than contextualizing it. As previously described, from our literature review, assessment, and subsequent interpretation of LoC scenarios, we derived that there exist no concrete LoC scenarios in the AI literature that describe an outcome below the national risk assessment threshold that would capture these types of events. Instead, we start to see concrete scenarios emerging around the middle of our graph (see Figure 2). [^20] We note that neither of the two previously mentioned examples, concerning OpenAI’s agent Operator and Replit, would trigger the national risk assessment threshold. Interpreting the LoC definition so broadly as to include similar instances would therefore ignore the reference points offered by AI literature, which rather suggest that the LoC definition was meant to capture a more significant dimension of losing the ability to “reliably direct” AI systems than instances of Deviation. Second, including instances of Deviation within the scope of the COP’s definition of LoC would ignore the qualification of LoC as a ‘systemic risk’ in the COP ([EU_AI_COP], Safety and Security Chapter, Appendix 1.4). Under the COP, systemic risks share the following “essential characteristics,” which speak to the severity and persistence of the LoC outcome ([EU_AI_COP], Safety and Security Chapter, Appendix 1.2.1): a “significant impact on the Union market” that “can be propagated at scale across the value chain.” Scenarios contained within Deviation and instances like the ones described above could hardly generate such a negative impact on the Union market. We therefore conclude that, while a literal interpretation of the definition offered by the COP could, in theory, include scenarios below the national risk assessment threshold (such as instances contained within Deviation), that was likely not its intent.
@@ 1.B.3 Strict Loss of Control
This category captures scenarios that cause absolute harm of the type that no action at any given point in the future would enable society to recover from the harmful event. The category is strict in the sense that it severely affects and is consequential to humanity as a whole and is permanent, and as such would be mapped at the top right corner of our graph (see Figure 2). “Existential catastrophes” are scenarios that involve the destruction of humanity’s long-term potential [ord2020precipice, sundaram_mani_2025, stauffer2023existential], for example, a nuclear winter that precipitates a prolonged agricultural collapse [sagan1983nuclear]. As such, the concept of existential catastrophe naturally delineates scenarios that are unusually serious and permanent, such as human extinction, from those that are ultimately reversible; and hence, acts as a natural boundary for separating Strict LoC scenarios from other, less absolute, forms of LoC (i.e. Bounded LoC).
@@ 1.C Reflections
We reached our LoC taxonomy by starting from a review of definitions (i) in AI literature and (ii) in other safety-critical sectors, including defense, aviation, and nuclear. We found that neither set of existing definitions supports one common, clear definition of LoC that could enable decision- and policymakers to operationalize LoC today. This result led us to, instead, home in on two leading multi-stakeholder-developed definitions of LoC, offered by the COP and the IASR. We observed that these two definitions remain difficult to operationalize, as they can encompass a wide range of outcomes and come into effect at different times. In order to shed more light on these definitions and on LoC as a concept more broadly, we subsequently conducted a literature review, which resulted in the assessment of 40 LoC scenarios. Specifically, we aimed to extrapolate the types of scenarios and consequent outcomes scholars are mostly concerned with. We developed a methodology to assess the scenarios, categorize them, and plot them on an experimental graph (see Figure 2).
@@ 2.A Capabilities and Propensities: Intuitive but not Actionable Today
Existing approaches to address the risk of LoC have concentrated on capturing and intervening on the appropriate AI capabilities [^22] An AI system’s capabilities refers to the behaviors the AI system can perform under ideal conditions. That is, it refers to the abilities the AI system has when it is optimally elicited and sufficiently resourced—which may be very different from the ability the AI system has in realistic conditions [Sharkey2024d]. and propensities [^23] An AI system’s propensities refers to the behaviors the AI system tends to display in real-world deployment conditions—for example, under real world prompting and safety mechanisms [Sharkey2024d]. that could trigger it. This focus is evident in both the COP and IASR texts, supplementing their LoC definitions [IASR_2025, EU_AI_COP]. However, decision- and policymakers seeking to identify and mitigate LoC threats today may find this challenging. There is no clear consensus on three points: (i) the specific capabilities and propensities that are necessary for an AI system to be able to cause LoC; (ii) the sub-capabilities within these capabilities; and (ii) the sub-capabilities that constitute these capabilities; and (iii) the critical thresholds after which specific capabilities could lead to LoC. In turn, this lack of consensus affects decision- and policymakers ability to suggest and implement adequate safeguards. We briefly present each of these three challenges in more depth below, before presenting a different approach that decision- and policymakers could operationalize today.

### v2

@@
The Loss of Control Playbook:
Degrees, Dynamics, and Preparedness
@@ Introduction
In turn, this allowed us to extrapolate some defining elements of LoC scenarios and to support decision- and policymakers in carving out the content and boundaries of LoC more clearly. Our analysis further enabled us to assess why LoC definitions across the literature may differ and to propose a novel conceptualization of LoC. Specifically, in our literature review, we found that LoC scenarios often differ in terms of their severity and persistence. These two metrics broadly capture intuitions around, among other things, the number of people affected, the degree to which those people are affected, and the difficulty of stopping the LoC incident. We subsequently used these metrics to plot concrete LoC scenarios from the literature (see footnote 15) onto an experimental graph, allowing us to provide additional context to existing descriptions of LoC.
@@ 1.A.1 Findings from Loss of Control Definitions in Literature
Despite certain commonalities, such as concerns about AI systems operating beyond the scope of reliable human direction or oversight, definitions in the broader AI literature differ in their emphasis and framing [bernardi2025societaladaptationadvancedai, kulveit2025gradualdisempowermentsystemicexistential, UK_DSIT_FrontierAI_2023, CIGI_Framework_Convention_2024] [^3] Due to our methodology, as described later in this Chapter, human disempowerment, due to its uncertain nature, is out of the scope of this report. thereby presenting conceptual distinctions. Divergences between definitions can make it difficult to implement functional frameworks and enact suitable interventions. One axis of variation relates to the cognitive capabilities required by an AI system for LoC. Some researchers appear to suggest LoC may occur once AI systems become more intelligent than humans [hendrycks2023overviewcatastrophicairisks, MIRI_Mission_Strategy_Update_2024]. Others more explicitly refer to artificial superintelligence when describing LoC risk [RAND_Rise_Fall_Nations, barnett2025aigovernanceavoidextinction]. In contrast, Stuart Russell’s definition in ‘Artificial Intelligence and the Problem of Control’ requires only a “sufficiently capable machine” [Russell2022], rather than one explicitly surpassing human intelligence or achieving superintelligence. A further divergence across definitions involves the difference in treatment between oversight and control. While some definitions refer to control [CA_Frontier_AI_Policy_2025, IDAIS_Shanghai_2025], others emphasize oversight instead, focusing on AI systems that “operat[e] beyond human oversight” [METR_AGI_Definitions_2025]. Another axis of variation concerns the possibility of regaining control. For example, the COP [EU_AI_COP] does not necessarily imply that LoC is irreversible, whereas the IASR [IASR_2025], the Singapore Consensus [bengio2025singaporeconsensusglobalai], and other reports [Bengio_managing_risks_2024, hendrycks2023overviewcatastrophicairisks] all suggest that LoC involves the absence of a clear path to regaining control. These interpretations imply that once control is lost, recovery may be impossible or extremely difficult. Overall, we found that while definitions aim to cover similar areas, they do so in sufficiently diverse ways to prevent a common, clear definition of LoC arising from the AI literature.
@@ 1.A.2 Findings from Loss of Control Scenarios in Literature
In total, we found 12 text passages that fulfilled all of our criteria and were therefore classified as concrete LoC scenarios (see Figure 1 for a high-level oversight of how our methodology worked).
@@
Next, we plotted all 12 concrete LoC scenarios on an experimental graph, using severity and persistence as the axes, [^13] Given the difficulty of computing the precise economic impact for persistence, especially for concrete LoC scenarios from the literature that do not tend to consider this aspect in any detail, we decided to use the same economic impact estimate for persistence as for severity. We note that, while severity and persistence are not necessarily equivalent in terms of economic impact in all instances, we believe that they are quasi-equivalent for the scenarios plotted on our graph (see footnote 15). and color-coded them within their respective threat categories (see footnote 15). In doing so, we noted that certain areas emerged from our plots and sought to contextualize them within concepts and frameworks that already exist in governance and which decision- and policymakers may be familiar with.
@@
First, we sought to contextualize the area indicated in the lower left of our graph (see footnote 15). In doing so, we came across the economic consequence thresholds that the U.S. Department of Homeland Security (DHS), alongside federal agencies and the intelligence community, identified as “necessary to create a national level-event” in the Strategic National Risk Assessment (SNRA) [DHS_SNRA_2011]. Specifically, in 2011, DHS led an effort to “identify the types of incidents that pose the greatest threat to the Nation’s homeland security,” including various natural, technological, and adversarial, human-caused hazards [DHS_SNRA_2011]. [^16] We decided to adopt the SNRA as a term of reference because the U.S. is the default location where the majority of literature assumes LoC scenarios to occur. For instance, ‘Scenario 1’ in [Kalra_GeopoliticsAGI] is an example of a scenario that envisages AI-caused electricity blackouts starting in the U.S. In some cases, the location of the LoC scenario described by literature is ambiguous (for instance, ‘Scenario 4’ in [Kalra_GeopoliticsAGI]), so for the purposes of making a comparable economic impact estimate, we assumed that the scenario occurred within the U.S. to ensure consistency across scenarios. For example, earthquakes, floods, hurricanes, wildfires, and cyberattacks against physical infrastructure meet the threshold if they result in direct economic losses exceeding \$100 million [DHS_SNRA_2011]. Similarly, a cyberattack against data [^17] In this context, data refers to the information contained in a computer system or data processes. is considered a national-level event if it results in economic losses of \$1 billion or more. In this research report, we adopt this last threshold set by the SNRA at \$1 billion or greater (adjusting it for inflation to approximately \$1.4 billion) to provide some contextualization of LoC scenarios (see footnote 15). We refer to this threshold, i.e., the \$1.4 billion, for convenience, as the ‘national risk assessment threshold.’ In summary, the national risk assessment threshold is the highest threshold based on economic consequences that DHS, federal agencies, and the intelligence community identified in the Strategic National Risk Assessment to demarcate a national-level event (i.e., a threat or hazard that has the potential to significantly impact the U.S. homeland security). We indicate this threshold in orange in Figure 2.
@@
Second, we sought to contextualize the area indicated in the top right of our graph (see footnote 15). In doing so, we drew on existing conceptualizations describing events of absolute scale and permanence. In other words, events that could lead to the destruction of humanity’s long-term potential [ord2020precipice, sundaram_mani_2025, stauffer2023existential, DallasFed_AI_Productivity_2025]. [^18] We note that we were unable to find a numerical operationalization for ‘existential catastrophe’ in the literature. Therefore, the boundary given on this graph is for illustrative purposes only, but is likely accurate, as it captures an absolute scale and permanence akin to the best interpretation of an existential catastrophe (see footnote 15). These events are commonly referred to as an ‘existential catastrophe,’ and we adopted this term to establish our upper boundary. In summary, existential catastrophe demarcates the point at which humanity loses control over its future in an absolute sense. We indicate this threshold in bold red dashes in Figure 2.
@@
  • Deviation: captures events that cause some harm or inconvenience, but lacks the requisite severity and persistence for inclusion within the threats and hazards that have the potential to significantly impact national preparedness. The events in this category would be plotted in the lower left corner of our graph (see footnote 15), and delineated from LoC scenarios in other categories through the national risk assessment threshold.
  • Bounded LoC: captures events that cause great damage or suffering, and are difficult, but possible, to contain, albeit potentially at great cost. The events in this category are mapped between the lower left corner and the top right corner.
  • Strict LoC: captures events which are maximally severe and permanent, such as events that result in humanity as a whole becoming extinct. The events in this category are mapped at the top right corner of our graph (see footnote 15).
@@ 1.B.1 Deviation
This category captures scenarios where an AI system briefly deviates from human intent, and, while causing some real-world damage or minor economic impacts, remains below the threshold for inclusion within the threats and hazards that have the potential to significantly impact the U.S. homeland security based on the Strategic National Risk Assessment [DHS_SNRA_2011]. These deviations from human intent are simple to stop at low or no cost, meaning that scenarios in this category would be plotted in the lower left corner on our graph (see footnote 15), and are delineated from other categories by what we term the national risk assessment threshold. Overall, we propose that the national risk assessment threshold is a principled boundary since it is calibrated to exclude events where harm is limited and containment is easy, that is, deviations.
@@
Nevertheless, we offer our reasoning as to why we do not believe the COP definition [EU_AI_COP] should be interpreted so broadly as to include the category of Deviation, derived from our earlier review and methodology. A broad interpretation of “losing the ability to reliably direct” would already capture scenarios that occur today. Consider the following two examples. In one instance, an AI system recently deleted an entire database, even though it was instructed not to modify the code [Cybernews_Replit_VibeCoding_2025]. In another instance, OpenAI’s agent Operator conducted a purchase without requesting the users’ consent, despite having been told only to find the cheapest option [Fowler_WaPo_Operator_2025]. This occurred even though the agent was trained to ask the humans’ consent before conducting irreversible actions such as purchasing goods or sending emails [OpenAI_introducing_operator]. In both cases, humans were not able to “reliably direct” an AI system and, therefore, both cases could, in theory, qualify as LoC based on a broad interpretation of “losing the ability to reliably direct.” However, we propose that it is unlikely the LoC definition in the COP was supposed to be interpreted so extensively as to include similar cases, which we believe qualify as Deviation rather than LoC. First, including the category of Deviation within the definition of LoC in the COP [EU_AI_COP] would run the risk of making the LoC definition overly broad, rather than contextualizing it. As previously described, from our literature review, assessment, and subsequent interpretation of LoC scenarios, we derived that there exist no concrete LoC scenarios in the AI literature that describe an outcome below the national risk assessment threshold that would capture these types of events. Instead, we start to see concrete scenarios emerging around the middle of our graph (see footnote 15). [^20] We note that neither of the two previously mentioned examples, concerning OpenAI’s agent Operator and Replit, would trigger the national risk assessment threshold. Interpreting the LoC definition so broadly as to include similar instances would therefore ignore the reference points offered by AI literature, which rather suggest that the LoC definition was meant to capture a more significant dimension of losing the ability to “reliably direct” AI systems than instances of Deviation. Second, including instances of Deviation within the scope of the COP’s definition of LoC would ignore the qualification of LoC as a ‘systemic risk’ in the COP ([EU_AI_COP], Safety and Security Chapter, Appendix 1.4). Under the COP, systemic risks share the following “essential characteristics,” which speak to the severity and persistence of the LoC outcome ([EU_AI_COP], Safety and Security Chapter, Appendix 1.2.1): a “significant impact on the Union market” that “can be propagated at scale across the value chain.” Scenarios contained within Deviation and instances like the ones described above could hardly generate such a negative impact on the Union market. We therefore conclude that, while a literal interpretation of the definition offered by the COP could, in theory, include scenarios below the national risk assessment threshold (such as instances contained within Deviation), that was likely not its intent.
@@ 1.B.3 Strict Loss of Control
This category captures scenarios that cause absolute harm of the type that no action at any given point in the future would enable society to recover from the harmful event. The category is strict in the sense that it severely affects and is consequential to humanity as a whole and is permanent, and as such would be mapped at the top right corner of our graph (see footnote 15). “Existential catastrophes” are scenarios that involve the destruction of humanity’s long-term potential [ord2020precipice, sundaram_mani_2025, stauffer2023existential], for example, a nuclear winter that precipitates a prolonged agricultural collapse [sagan1983nuclear]. As such, the concept of existential catastrophe naturally delineates scenarios that are unusually serious and permanent, such as human extinction, from those that are ultimately reversible; and hence, acts as a natural boundary for separating Strict LoC scenarios from other, less absolute, forms of LoC (i.e. Bounded LoC).
@@ 1.C Reflections
We reached our LoC taxonomy by starting from a review of definitions (i) in AI literature and (ii) in other safety-critical sectors, including defense, aviation, and nuclear. We found that neither set of existing definitions supports one common, clear definition of LoC that could enable decision- and policymakers to operationalize LoC today. This result led us to, instead, home in on two leading multi-stakeholder-developed definitions of LoC, offered by the COP and the IASR. We observed that these two definitions remain difficult to operationalize, as they can encompass a wide range of outcomes and come into effect at different times. In order to shed more light on these definitions and on LoC as a concept more broadly, we subsequently conducted a literature review, which resulted in the assessment of 40 LoC scenarios. Specifically, we aimed to extrapolate the types of scenarios and consequent outcomes scholars are mostly concerned with. We developed a methodology to assess the scenarios, categorize them, and plot them on an experimental graph (see footnote 15).
@@ 2.A Capabilities and Propensities: Intuitive but not Actionable Today
Existing approaches to address the risk of LoC have concentrated on capturing and intervening on the appropriate AI capabilities [^22] An AI system’s capabilities refers to the behaviors the AI system can perform under ideal conditions. That is, it refers to the abilities the AI system has when it is optimally elicited and sufficiently resourced—which may be very different from the ability the AI system has in realistic conditions [Sharkey2024d]. and propensities [^23] An AI system’s propensities refers to the behaviors the AI system tends to display in real-world deployment conditions—for example, under real world prompting and safety mechanisms [Sharkey2024d]. that could trigger it. This focus is evident in both the COP and IASR texts, supplementing their LoC definitions [IASR_2025, EU_AI_COP]. However, decision- and policymakers seeking to identify and mitigate LoC threats today may find this challenging. There is no clear consensus on three points: (i) the specific capabilities and propensities that are necessary for an AI system to be able to cause LoC; (ii) the sub-capabilities that constitute these capabilities; and (iii) the critical thresholds after which specific capabilities could lead to LoC. In turn, this lack of consensus affects decision- and policymakers’ ability to suggest and implement adequate safeguards. We briefly present each of these three challenges in more depth below, before presenting a different approach that decision- and policymakers could operationalize today.